OpenStudy (anonymous):

Is it possible for an equation to have no verticle, oblique or horizontal retricemptotes?

OpenStudy (anonymous):

you mean asymptotes? yes, examples y=x^2 y=mx+b

OpenStudy (anonymous):

if the 2 terms mean the same thing then my response is valid.